# Commutation relations for two-sided radial SLE

### v2

@@ Abstract
We study the commutation relation for 2-radial SLE in the unit disc starting from two boundary points. We follow the framework introduced by Dubédat [9]. Under an additional requirement of the interchangeability of the two curves, we classify all locally commuting 2-radial SLEκ for $\kappa\in(0,8)$: it is either a two-sided radial SLEκ with spiral of constant spiraling rate or a chordal SLEκ weighted by a power of the conformal radius of its complement. Namely, for fixed $\kappa$ and starting points, we have exactly two one-parameter continuous families of locally commuting 2-radial SLE. Two-sided radial SLE with spiral is a generalization of two-sided radial SLE (without spiral) analyzed in [20, 13, 23, 14] and satisfies the resampling property. We also discuss the semiclassical limit of the commutation relation as $\kappa\to 0$. In particular, we show that the limit for the second family with an appropriately chosen power of conformal radius is a chord that minimizes a modified chordal Loewner energy, which is unique only when the endpoints are not antipodal.
Keywords: Schramm–Loewner evolution, commutation relation, resampling property.
@@ 1.1 Background on radial SLE
However, to characterize the other natural variant — radial SLE — one needs an additional condition on the reflection symmetry. As we are mainly concerned with radial SLE in the present article, let us briefly describe its definition and characterization. We will describe the radial Loewner chain in $\mathbb{D}=\{z\in\mathbb{C}\,|\,|z|<1\}$ targeting at 00. Radial Loewner chain in other simply connected domain $D\subset\mathbb{C}$ targeting at an interior point $z_{0}\in D$ is defined via a conformal map from $\mathbb{D}$ onto $D$ sending 00 to $z_{0}$.
Conformal radius and capacity. For any compact subset $K$ (not necessarily connected) of $\overline{\mathbb{D}}$ such that $\mathbb{D}\setminus K$ is simply connected and contains 00, we let $g_{K}$ be the unique conformal map $\mathbb{D}\setminus K\to\mathbb{D}$ such that $g_{K}(0)=0$ and $g^{\prime}_{K}(0)>0$ (called the radial mapping-out function of $K$). The conformal radius of $\mathbb{D}\setminus K$ is
@@
Radial Loewner chain. For $\theta\in[0,2\pi)$, suppose $\eta:[0,T]\to\overline{\mathbb{D}}$ is a continuous non-self-crossing curve such that $\eta_{0}=\mathrm{e}^{\mathfrak{i}\theta}$ and $\eta_{(0,T)}\subset\mathbb{D}\setminus\{0\}$. Let $D_{t}$ be the connected component of $\mathbb{D}\setminus\eta_{[0,t]}$ containing the origin. Let $g_{t}:D_{t}\to\mathbb{D}$ be the unique conformal map with $g_{t}(0)=0$ and $g_{t}^{\prime}(0)>0$. We say that the curve is parameterized by capacity if $g_{t}^{\prime}(0)=\exp(t)$. Then $g_{t}$ satisfies the radial Loewner equation
@@
Characterization of radial SLE. Consider a family $(\mathbb{P}_{\theta})_{\theta\in\mathbb{R}/2\pi\mathbb{Z}}$ of probability distributions on curves $\eta:[0,\infty)\to\overline{\mathbb{D}}$ with $\eta_{0}=\mathrm{e}^{\mathfrak{i}\theta}$ and parametrized by capacity satisfies the following properties:
@@
    (Conformal invariance) For all $a\in\mathbb{R}/2\pi\mathbb{Z}$, let $\rho_{a}(z)=\mathrm{e}^{\mathfrak{i}a}z$ be the rotation map $\mathbb{D}\to\mathbb{D}$. For all $a,\theta\in\mathbb{R}/2\pi\mathbb{Z}$, the pullback measure $\rho_{a}^{*}\,\mathbb{P}_{\theta}=\mathbb{P}_{\theta-a}$. From this, we may extend the definition of $\mathbb{P}_{\theta}$ to any simply connected domain $D$ with an interior marked point $z_{0}$ by pulling back using a uniformizing conformal map $D\to\mathbb{D}$ sending $z_{0}$ to 00.
@@ 1.2 Locally commuting 2-radial SLE
The random radial curve satisfying conformal invariance and domain Markov property are easily characterized thanks to the fact all simply connected domains (excluding $\mathbb{C}$) with one marked interior point and one marked boundary point are conformally equivalent. If we consider the radial Loewner chain of two curves growing from two distinct boundary points $\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}}\in\partial\mathbb{D}$, then we have a one-dimensional moduli space of the boundary data. The goal of this work is to characterize and identify the family of local laws $\mathbb{P}_{(\theta_{1},\theta_{2})}$ of the pair of continuous non-self-crossing curves $(\eta^{(1)},\eta^{(2)})$ in $\mathbb{D}$ starting from all possible choices of $\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}}\in\partial\mathbb{D}$ under the following axioms.
For this, we first parameterize $\eta^{(1)},\eta^{(2)}$ by their intrinsic radial capacities and consider the associated Loewner chains $g_{\cdot}^{(1)}$ and $g_{\cdot}^{(2)}$ respectively. For all $\mathbf{t}=(t_{1},t_{2})\in\mathbb{R}_{>0}^{2}$, let $g_{\mathbf{t}}$ be the radial mapping-out function of $\eta_{\mathbf{t}}:=\eta^{(1)}_{[0,t_{1}]}\cup\eta^{(2)}_{[0,t_{2}]}$.
@@
    Conformal invariance: For all $a\in\mathbb{R}/2\pi\mathbb{Z}$ and $\theta_{1},\theta_{2}\in\mathbb{R}/2\pi\mathbb{Z}$, the pullback measure $\rho_{a}^{*}\,\mathbb{P}_{(\theta_{1},\theta_{2})}=\mathbb{P}_{(\theta_{1}-a,\theta_{2}-a)}$. From this, we may extend the definition of $\mathbb{P}_{\theta_{1},\theta_{2}}$ to any simply connected domain $D$ with an interior marked point $z_{0}$ by pulling back using a uniformizing conformal map $D\to\mathbb{D}$ sending $z_{0}$ to 00.
@@
    where $(\exp(\mathfrak{i}\theta^{(1)}_{\mathbf{t}}),\exp(\mathfrak{i}\theta^{(2)}_{\mathbf{t}}))=(g_{\mathbf{t}}(\eta^{(1)}_{t_{1}}),g_{\mathbf{t}}(\eta^{(2)}_{t_{2}}))$.
@@
See Section 2.1 for a more precise statement of the axioms. We note that since $g_{\mathbf{t}}$ is parametrized by two times $t_{1},t_{2}$, the condition (DMP) implies that $g_{\mathbf{t}}$ may be realized by first mapping out $\eta^{(1)}_{[0,t_{1}]}$ using $g_{t_{1}}^{(1)}$, then mapping out $g_{t_{1}}^{(1)}(\eta^{(2)}_{[0,t_{2}]})$, or vice versa. The image has the same law regardless of the order in which we map out the curves. This observation gives a commutation relation on the infinitesimal generators of the two curves (Proposition 2.2). Therefore, we call the family of the laws $(\mathbb{P}_{(\theta_{1},\theta_{2})})$ as an interchangeable and locally commuting 2-radial SLEκ.
These conditions and our analysis are very close to the commutation relation of SLE studied by Dubédat in [9], which focuses on classifying all locally commuting chordal SLEs. Dubédat also derived the commutation relation in the radial setting in terms of the generators of radial SLE. Our main contribution is to find all solutions to the radial commutation relation and identify all interchangeable and locally commuting 2-radial SLEκ. It is also natural to use SLE/GFF couplings, particularly, [28], to find examples of commuting $2$-radial SLEs. See Remark 3.10. However, it is unclear to the authors a priori that all commuting $2$-radial SLEs can be coupled to GFF.
@@ 1.3 Main result: A classification
Both $\mathcal{G}_{\mu}$ in (1.3) and $\mathcal{Z}_{\alpha}$ in (1.4) are solutions to “radial BPZ equations” (or “BPZ-Cardy equations” in [18], see also [19]):
@@
$F$ $=\frac{\mu^{2}-3}{2\kappa}\geq-\frac{3}{2\kappa},\quad$ when $\text{when }\mathcal{Z}=\mathcal{G}_{\mu},$
$F$ $=\frac{(6-\kappa)(\kappa-2)}{8\kappa}-\alpha>-\frac{3}{2\kappa},\quad$ when $\text{when }\mathcal{Z}=\mathcal{Z}_{\alpha}.$
In general, partition function $\mathcal{Z}_{\alpha}$ in (1.4) involves hypergeometric function and is not explicit; but in the following two special cases, such function has a simple explicit expression:
@@
  • When $\alpha=\alpha_{1}(\kappa)$ which is the one-arm exponent for the conformal loop ensemble, $\mathcal{Z}_{\alpha_{1}(\kappa)}$ can be written as trigonometric functions in $\theta=\theta_{2}-\theta_{1}$, see Remark 2.7. In this case, the partition function $\mathcal{Z}_{\alpha_{1}(\kappa)}$ with $\kappa\in(4,8)$ is the limit of the partition function for interfaces in critical random-cluster model conditional on the one-arm event.
The two-sided radial SLEκ with spiral is also closely related to the mixed multifractal spectrum introduced by Binder [5]. More precisely, the points on an SLE curve are classified in terms of the asymptotic behavior of the uniformizing conformal maps in the complement of the curve, according to its Hölder exponent and its winding exponent. The Hausdorff dimension (i.e., mixed multifractal spectrum) of the set of points with a given behavior is computed in [11]. It was shown in [23] that conditioning a chordal SLE to pass through an interior point 00 gives the two-sided radial SLE (without spiral). Since the points with zero winding exponent have the maximal spectrum, this conditioning is equivalent to conditioning on the event where the 00 is a point on the SLE curve with zero winding exponent. We speculate that “conditioning” on the rare event where 00 is a point with winding rate $\mu$, the curve obtained should satisfy the commutation relation and hence has to be the two-sided radial SLE that with the spiraling rate $\mu$. In the terminology of the conformal field theory, this would correspond to inserting a curve-generating operator at 00 with complex charges and conformal weights [4].
@@ 1.4 Resampling property
Let $D\subset\mathbb{C}$ be a simply connected domain and let $x_{1},x_{2}$ be distinct prime ends of $\partial D$. We denote by $\mathfrak{X}(D;x_{1},x_{2})$ the set of continuous non-self-crossing curves in $D$ from $x_{1}$ to $x_{2}$. We recall the definition of the chordal SLE in the upper half-plane $({\mathbb{H}};0,\infty)$ in Appendix B. For other simply connected domain $(D;x_{1},x_{2})$, chordal SLE in $(D;x_{1},x_{2})$ is a random curve in $\mathfrak{X}(D;x_{1},x_{2})$ defined by mapping chordal SLE in $({\mathbb{H}};0,\infty)$ conformally from ${\mathbb{H}}$ onto $D$ sending 00 to $x_{1}$ and $\infty$ to $x_{2}$. Let $D\subset\mathbb{D}$ be a simply connected domain containing 00 and let $x_{1},x_{2}$ be distinct prime ends of $\partial D$. We denote by $\mathfrak{X}(D;x_{1},x_{2};0)$ the set of pairs of continuous simple curves $(\eta^{(1)},\eta^{(2)})$ in $D$ such that $\eta^{(1)}$ goes from $x_{1}$ to 00 and $\eta^{(2)}$ goes from $x_{2}$ to 00 and $\eta^{(1)}\cap\eta^{(2)}=\{0\}$. We say that a law on $(\eta^{(1)},\eta^{(2)})\in\mathfrak{X}(D;x_{1},x_{2};0)$ satisfies resampling property if
@@
Fix $\kappa\in(0,4]$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. An interchangeable and locally commuting 2-radial $\operatorname{SLE}_{\kappa}$ in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ is a two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral, hence, satisfies the resampling property. The converse is not true as we may take a linear combination of 2-radial $\operatorname{SLE}_{\kappa}$ with different spiraling rate.
@@ 1.5 Semiclassical limits of partition functions
We also discuss the commutation relation when $\kappa=0$, the corresponding deterministic pair of curves, and the semiclassical limit $\kappa\to 0{\scriptstyle+}$ in Section 4. In particular, semiclassical limits of partition functions in Theorem 1.1 have explicit formulas.
@@
    $\lim_{\kappa\to 0}\kappa\log\mathcal{G}_{\mu}(\theta_{1},\theta_{2})=2\log\sin\left((\theta_{2}-\theta_{1})/2\right)+\mu(\theta_{1}+\theta_{2}).$ (1.6)
@@
We mention that a similar semiclassical limit of the partition functions for multichordal SLE was considered in [29, 3] and that multiradial $\operatorname{SLE}_{0}$ was considered in [2, 40]. We prove Proposition 1.3 in Section 4.

### v3

@@ Abstract
We study the commutation relation for 2-radial SLE in the unit disc starting from two boundary points. We follow the framework introduced by Dubédat [12]. Under an additional requirement of the interchangeability of the two curves, we classify all locally commuting 2-radial SLEκ for $\kappa\in(0,8)$: it is either a two-sided radial SLEκ with spiral of constant spiraling rate or a chordal SLEκ weighted by a power of the conformal radius of its complement. Namely, for fixed $\kappa$ and starting points, we have exactly two one-parameter continuous families of locally commuting 2-radial SLE. We also discuss the semiclassical limit of the commutation relation as $\kappa\to 0$. In particular, we show that the limit for the second family with an appropriately chosen power of conformal radius is a chord that minimizes a modified chordal Loewner energy, which is unique only when the endpoints are not antipodal.
Keywords: Schramm–Loewner evolution, radial Loewner chain, commutation relation, resampling property.
@@ 1.1 Background on radial SLE
However, to characterize the other natural variant — radial SLE — one needs an additional condition on the reflection symmetry. As we are mainly concerned with radial SLE in the present article, let us briefly describe its definition and characterization. We will describe the radial Loewner chain in $\mathbb{D}=\{z\in\mathbb{C}\,|\,|z|<1\}$ targeting at 0. Radial Loewner chain in other simply connected domain $D\subset\mathbb{C}$ targeting at an interior point $z_{0}\in D$ is defined via a conformal map from $\mathbb{D}$ onto $D$ sending 0 to $z_{0}$.
Conformal radius and capacity. For any compact subset $K$ (not necessarily connected) of $\overline{\mathbb{D}}$ such that $\mathbb{D}\setminus K$ is simply connected and contains 0, we let $g_{K}$ be the unique conformal map $\mathbb{D}\setminus K\to\mathbb{D}$ such that $g_{K}(0)=0$ and $g^{\prime}_{K}(0)>0$ (called the radial mapping-out function of $K$). The conformal radius of $\mathbb{D}\setminus K$ is
@@
Radial Loewner chain. For $\theta\in[0,2\pi)$, suppose $\eta:[0,T]\to\overline{\mathbb{D}}$ is a continuous non-self-crossing curve such that $\eta_{0}=\mathrm{e}^{\mathfrak{i}\theta}$ and $\eta_{(0,T)}\subset\mathbb{D}\setminus\{0\}$. Let $D_{t}$ be the connected component of $\mathbb{D}\setminus\eta_{[0,t]}$ containing the origin. Let $g_{t}:D_{t}\to\mathbb{D}$ be the unique conformal map with $g_{t}(0)=0$ and $g_{t}^{\prime}(0)>0$. We say that the curve is parameterized by capacity if $g_{t}^{\prime}(0)=\mathrm{e}^{t}$. Then $g_{t}$ satisfies the radial Loewner equation
@@
Characterization of radial SLE. Consider a family $(\mathbb{P}_{\theta})_{\theta\in\mathbb{R}/2\pi\mathbb{Z}}$ of probability distributions on curves $\eta:[0,\infty)\to\overline{\mathbb{D}}$ with $\eta_{0}=\mathrm{e}^{\mathfrak{i}\theta}$ and parametrized by capacity that satisfies the following properties:
@@
    (Conformal invariance) For all $a\in\mathbb{R}/2\pi\mathbb{Z}$, let $\rho_{a}(z)=\mathrm{e}^{\mathfrak{i}a}z$ be the rotation map $\mathbb{D}\to\mathbb{D}$. For all $a,\theta\in\mathbb{R}/2\pi\mathbb{Z}$, the pullback measure $\rho_{a}^{*}\,\mathbb{P}_{\theta}=\mathbb{P}_{\theta-a}$. From this, we may extend the definition of $\mathbb{P}_{\theta}$ to any simply connected domain $D$ with an interior marked point $z_{0}$ by pulling back using a uniformizing conformal map $D\to\mathbb{D}$ sending $z_{0}$ to 0.
@@ 1.2 Locally commuting 2-radial SLE
The random radial curve satisfying conformal invariance and domain Markov property are easily characterized thanks to the fact all simply connected domains (excluding $\mathbb{C}$) with one marked interior point and one marked boundary point are conformally equivalent. If we consider the radial Loewner chain of two curves growing from two distinct boundary points $\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}}\in\partial\mathbb{D}$, then we have a one-dimensional moduli space of the boundary data that we need to take into account.
The goal of this work is to give an axiomatic characterization of locally commuting 2-radial SLEs. One example of such curves of particular interest, is the two-sided radial SLE analyzed in [25, 16, 28, 17] which arises as the chordal SLE in $\mathbb{D}$ connecting two boundary points conditioned on passing through 0.
More precisely, locally commuting 2-radial SLE is a family of local laws $\mathbb{P}_{(\theta_{1},\theta_{2})}$ of the pair of continuous non-self-crossing curves $(\eta^{(1)},\eta^{(2)})$ in $\mathbb{D}$ starting from all possible choices of $\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}}\in\partial\mathbb{D}$ under the following axioms.
We first parameterize $\eta^{(1)},\eta^{(2)}$ by their intrinsic radial capacities and consider the associated radial Loewner chains $g_{\cdot}^{(1)}$ and $g_{\cdot}^{(2)}$ respectively. For all $\mathbf{t}=(t_{1},t_{2})\in\mathbb{R}_{>0}^{2}$, let $g_{\mathbf{t}}$ be the radial mapping-out function of $\eta_{\mathbf{t}}:=\eta^{(1)}_{[0,t_{1}]}\cup\eta^{(2)}_{[0,t_{2}]}$.
@@
    Conformal invariance: For all $a\in\mathbb{R}/2\pi\mathbb{Z}$ and $\theta_{1},\theta_{2}\in\mathbb{R}/2\pi\mathbb{Z}$, the pullback measure $\rho_{a}^{*}\,\mathbb{P}_{(\theta_{1},\theta_{2})}=\mathbb{P}_{(\theta_{1}-a,\theta_{2}-a)}$. From this, we may extend the definition of $\mathbb{P}_{\theta_{1},\theta_{2}}$ to any simply connected domain $D$ with an interior marked point $z_{0}$ by pulling back using a uniformizing conformal map $D\to\mathbb{D}$ sending $z_{0}$ to 0.
@@
    where $\left(\exp(\mathfrak{i}\theta^{(1)}_{\mathbf{t}}),\exp(\mathfrak{i}\theta^{(2)}_{\mathbf{t}})\right)=\left(g_{\mathbf{t}}(\eta^{(1)}_{t_{1}}),g_{\mathbf{t}}(\eta^{(2)}_{t_{2}})\right)$.
@@
We note that, by “radial” we mean that the conformal mapping-out functions are normalized at the interior point $0\in\mathbb{D}$ and the precise axioms are written in neighborhoods of the starting points of the curves, see Section 2.1. In particular, they do not necessarily imply that the curves end at 0.
We also note that since $g_{\mathbf{t}}$ is parametrized by two times $t_{1},t_{2}$, the condition (DMP) implies that $g_{\mathbf{t}}$ may be realized by first mapping out $\eta^{(1)}_{[0,t_{1}]}$ using $g_{t_{1}}^{(1)}$, then mapping out $g_{t_{1}}^{(1)}(\eta^{(2)}_{[0,t_{2}]})$, or vice versa. The image has the same law regardless of the order in which we map out the curves. This observation gives a commutation relation on the infinitesimal generators of the two curves (Proposition 2.2). Therefore, we call the family of the laws $(\mathbb{P}_{(\theta_{1},\theta_{2})})$ as an interchangeable and locally commuting 2-radial SLEκ.
These conditions and our analysis are very close to the commutation relation of SLE studied by Dubédat in [12], which focuses on classifying all locally commuting chordal SLEs. Dubédat also derived the commutation relation in the radial setting in terms of the generators of radial SLE. One of our contributions is to find all solutions to the radial commutation relation and identify all interchangeable and locally commuting 2-radial SLEκ. It is also natural to use SLE/GFF couplings, particularly, [33], to find examples of commuting $2$-radial SLEs. See Remark 3.9. Nevertheless, our approach is to give an axiomatic characterization by studying the BPZ equations satisfied by the partition functions, and it is unclear to the authors that all commuting $2$-radial SLEs can be coupled to GFF a priori.
@@ 1.3 Main result: A classification
The first class above — two-sided radial SLE with spiral — also appeared in [33] in the context of SLE/GFF coupling. We will discuss other related literature in Section 1.6. The authors are unaware of studies on the second class of curves.
Both $\mathcal{G}_{\mu}$ in (1.3) and $\mathcal{Z}_{\alpha}$ in (1.4) are solutions to radial Belavin–Polyakov–Zamolodchikov (BPZ) equations (or called BPZ–Cardy equations derived in [23, 24], see also [4, 12] for the BPZ equation for multichordal SLE):
@@
$F$ $=\frac{\mu^{2}-3}{2\kappa}\geq-\frac{3}{2\kappa},\quad$ $\text{when }\mathcal{Z}=\mathcal{G}_{\mu},$
$F$ $=\frac{(6-\kappa)(\kappa-2)}{8\kappa}-\alpha>-\frac{3}{2\kappa},\quad$ $\text{when }\mathcal{Z}=\mathcal{Z}_{\alpha}.$
In general, the partition function $\mathcal{Z}_{\alpha}$ in (1.4) involves hypergeometric function and is not explicit; but in the following two special cases, such function has a simple explicit expression:
@@
  • When $\alpha=\alpha_{1}(\kappa)$ which is the one-arm exponent for the conformal loop ensemble, $\mathcal{Z}_{\alpha_{1}(\kappa)}$ can be written as trigonometric functions in $\theta=\theta_{2}-\theta_{1}$, see Remark 2.7.
@@ 1.4 Resampling property
Let $D\subset\mathbb{C}$ be a simply connected domain and let $x_{1},x_{2}$ be distinct prime ends of $\partial D$. We denote by $\mathfrak{X}(D;x_{1},x_{2})$ the set of continuous non-self-crossing curves in $D$ from $x_{1}$ to $x_{2}$. For other simply connected domain $(D;x_{1},x_{2})$, chordal SLE in $(D;x_{1},x_{2})$ is a random curve in $\mathfrak{X}(D;x_{1},x_{2})$ defined by mapping chordal SLE in $({\mathbb{H}};0,\infty)$ conformally from $\mathbb{H}$} onto $D$ sending 0 to $x_{1}$ and $\infty$ to $x_{2}$. Let $D\subset\mathbb{D}$ be a simply connected domain containing 0 and let $x_{1},x_{2}$ be distinct prime ends of $\partial D$. We denote by $\mathfrak{X}(D;x_{1},x_{2};0)$ the set of pairs of continuous simple curves $(\eta^{(1)},\eta^{(2)})$ in $D$ such that $\eta^{(1)}$ goes from $x_{1}$ to 0 and $\eta^{(2)}$ goes from $x_{2}$ to 0 and $\eta^{(1)}\cap\eta^{(2)}=\{0\}$. We say that a law on $(\eta^{(1)},\eta^{(2)})\in\mathfrak{X}(D;x_{1},x_{2};0)$ satisfies resampling property if
@@
Fix $\kappa\in(0,4]$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. If an interchangeable and locally commuting 2-radial $\operatorname{SLE}_{\kappa}$ is in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$, then it is a two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral, hence, it satisfies the resampling property. The converse is not true: a linear combination of two-sided radial $\operatorname{SLE}_{\kappa}$ with different spiraling rates is in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ and satisfies the resampling property but is not a locally commuting 2-radial $\operatorname{SLE}_{\kappa}$.
@@ 1.5 Semiclassical limits of partition functions
We also discuss the commutation relation when $\kappa=0$, the corresponding deterministic pair of curves, and the semiclassical limit $\kappa\to 0{\scriptstyle+}$ in Section 4. In particular, the semiclassical limits of partition functions in Theorem 1.1 have explicit formulas.
@@
    $\lim_{\kappa\to 0}\kappa\log\mathcal{G}_{\mu}(\theta_{1},\theta_{2})=2\log\sin\big((\theta_{2}-\theta_{1})/2\big)+\mu(\theta_{1}+\theta_{2}).$ (1.6)
@@
We prove Proposition 1.3 in Section 4. We mention that a similar semiclassical limit of the partition functions for multichordal SLE was considered in [34, 3] and that multiradial $\operatorname{SLE}_{0}$ was considered in [2, 45] (see also Section 1.6).
@@
1.6 Comments and further developments
Let us first make some comments on the related works to two-sided radial SLE with spiral. Apart from [33] mentioned above in the context of imaginary geometry, we believe it to be also related to the mixed multifractal spectrum introduced by Binder [6]. More precisely, the points on an SLE curve are classified in terms of the asymptotic behavior of the uniformizing conformal maps in the complement of the curve, according to its Hölder exponent and its winding exponent. The Hausdorff dimension (i.e., mixed multifractal spectrum) of the set of points with a given behavior is computed in [14]. It was shown in [28] that conditioning a chordal SLE to pass through an interior point 0 gives the two-sided radial SLE (without spiral). Since the points with zero winding exponent have the maximal spectrum, this conditioning is equivalent to conditioning on the event where the 0 is a point on the SLE curve with zero winding exponent. We speculate that “conditioning” on the rare event where 0 is a point with winding rate $\mu$, the curve obtained should satisfy the commutation relation and hence has to be the two-sided radial SLE that with the spiraling rate $\mu$. In the terminology of the conformal field theory, this would correspond to inserting a curve-generating operator at 0 with complex charges and conformal weights [5].
There are also a few recent works built upon this article. A follow-up paper [19] generalizes the first class in Theorem 1.1—two-sided radial SLE with spiral—to the multi-sided case and derives its resampling and boundary perturbation properties. Another follow-up paper [15] generalizes the second class in Theorem 1.1 to the multi-chordal case and connects the corresponding partition functions with $\alpha=\alpha_{1}(\kappa)$ to random-cluster model in polygon conditional on the one-arm event. Recent papers [45, 46] of J. Zhang continue the investigation of the radial commutation by studying locally commuting $n$-radial SLE and the corresponding radial BPZ equations.Zhang constructs solutions to these radial BPZ equations using Coulomb gas formalism and relates these solutions to quantum Calogero–Sutherland system in conformal field theory.
Regarding related works on semiclassical limits of SLEs, apart from those mentioned in Section 1.5, we also mention that the resampling property in the deterministic $\kappa=0$ case is equivalent to the geodesic property studied in [29, 22] (and the two-sided radial SLE0 is a geodesic pair in their terminology). See also [8, 34, 7] for other occurrences of geodesic property. Zhang also studied the $\kappa=0$ dynamics of multiple radial SLE0 [45] in relation to the classical Calogero–Sutherland system (see also [3] in the multichordal case).